OpenStudy (anonymous):

Help please :c Having survived the meteor impact, thanks to some last minute evasive maneuvers, the Mathonauts now set their sights on their Interstellar Headquarters. The Interstellar Headquarters orbits the Earth based on the equation y2 + x2 = 40,000. Using the original trajectory of the ship and complete sentences, explain to the pilot how to find where the ship’s path will cross the Interstellar Headquarters’s path. y=2x+1

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

So far I got 4x^4+4x+1=40,000

ganeshie8 (ganeshie8):

so you want to solve below equations to knw where the both trajectories intersect is it ? \( y^2 + x^2 = 40,000\) \(y=2x+1\)

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

which I put y into the equation and squared it getting 4x^2+4x+1+x^2=40,000

ganeshie8 (ganeshie8):

looks good^

ganeshie8 (ganeshie8):

get it into standard quadratic form, then apply quadratic formula

OpenStudy (anonymous):

so I have to set one side equal to zero?

OpenStudy (anonymous):

I'm sorry, I'm terrible at math.

ganeshie8 (ganeshie8):

\(4x^2+4x+1+x^2=40,000\) combine like terms \(5x^2+4x+1=40,000\) subtract 40,000 both sides \(5x^2+4x - 39,999= 0 \)

ganeshie8 (ganeshie8):

use quadratic formula and use ur calculator

OpenStudy (anonymous):

You're going to hate me for saying this but I have no clue where to start

ganeshie8 (ganeshie8):

lol no, below is the quadratic formula : \[\dfrac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

okay let me solve it, one second (:

ganeshie8 (ganeshie8):

your equation : \(5x^2+4x - 39,999= 0\) a = 5 b = 4 c = -39,999

ganeshie8 (ganeshie8):

plug them in the formula and USE UR CALCULATOR. u cannot solve it by hand

OpenStudy (anonymous):

Too much work by hand lol

ganeshie8 (ganeshie8):

it gives u two solutions, u take whichever u want..

ganeshie8 (ganeshie8):

or both... depending on which direction the ship/satellite are going

OpenStudy (anonymous):

i got \[-4\sqrt{-799980}\25\]

ganeshie8 (ganeshie8):

\[\dfrac{-4 \pm \sqrt{4^2-4(5)(-39,999)}}{2(5)}\] \[\dfrac{-4 \pm \sqrt{4^2+799980}}{10}\]

ganeshie8 (ganeshie8):

simplify further

ganeshie8 (ganeshie8):

\[\dfrac{-4 \pm \sqrt{4^2-4(5)(-39,999)}}{2(5)}\] \[\dfrac{-4 \pm \sqrt{4^2+799980}}{10}\] \[\dfrac{-4 \pm 894.4}{10}\]

OpenStudy (anonymous):

Okay, that wouldn't be the intersecting point though, right?

OpenStudy (anonymous):

Thank you so much!!

ganeshie8 (ganeshie8):

^^there ur both intersecting points A and B graphing is the best method to solve these equations i think

OpenStudy (anonymous):

Which one would I use though? point a or b?

OpenStudy (anonymous):

Or would I include both?

ganeshie8 (ganeshie8):

\(A = (-89.84, -178.68)\) \(B = (89.04, 179.08)\)

ganeshie8 (ganeshie8):

I think you only want the point B

OpenStudy (anonymous):

You're a life saver man!

OpenStudy (anonymous):

Thanks again

ganeshie8 (ganeshie8):

u wlc :)